An asteroid has a surface area of 5 m2. What would its surface area be if its linear dimension is doubled?

Please, explain to me how should I approach this problem because I really don't get it :-(
Answer #1

Assuming that asteroid is spherical in shape, its surface area A= 4*pi*r2

For a sphere , linear dimension is its diameter or radius

Where r is its radius. If r is doubled: new r =r' = 2r

Then Surface area = A' = 4 * pi * r'2 = 4 * pi * (2r)2 = 4 * 4*pi*r2 = 4 * A

or,

New surface area A' = 4 times old Surface area = 4A =4 * 5 = 20 m2
Answer #3

surface area=4 x pi x r2=pix(2r)2=pi x d2

A1=pi d12 and A2=pi x d22

A1/A2=(d1/d2)2

A2=(d2/d1)2 A1=(2d1/d1)2 x 5=20m2